### v1

@@ 4 Radial velocity analysis
For the radial velocity analysis we utilise Generalized Lomb Scargle (GLS) periodograms (Zechmeister & Kürster, 2009). To determine the false alarm probability levels (FAP) we used two methods. For peaks with significance lower than 10-5 and the 10%, 1%, 0.1% FAP levels in the periodograms, we used a bootstrap algorithm with 105 iterations. The method consists in generating a time series with dates similar to the real series of measurements, but assigning to each of them a RV measurement chosen randomly from the original data series. We label these FAP levels FAPbootstrap. For peaks with significance higher than 10-5 (i.e. inferior to 1/(number of boostrap iterations)), we used the GLS FAP analytical implementation of the package pyAstronomy (Czesla et al., 2019). We label these FAP levels FAPanalytical.
@@ 4.1 SOPHIE radial velocities
We have achieved with SOPHIE an average RV precision per measurement of 2.4 m s-1 and a rms of 13.9 m s-1 in the full time-series. Figure 2 displays the GLS periodograms obtained for the full data set, the 20102012 and the 20212023 periods respectively. Table 2 summarizes the properties of the time-series and the strongest peak in the periodogram.
The full data-set displays a peak at days with a FAPbootstrap of 4.0. This peak corresponds to the /2 harmonic of the stellar rotation period of days (Donati et al., 2023). Besides the 1 day alias there are no other peaks in the full data-set periodogram. No signal is observed at the stellar rotation period. The 20102012 season shows no significant periodicities. A broad peak is seen close to 6.99 days but its significance is low (FAPbootstrap=0.2). In the 20212023 period, the days appears at a FAPbootstrap level of 3.810-3. Besides the 1 day alias, no peaks are present in the periodogram with significance below a FAPbootstrap=10% level.
In summary, in the raw SOPHIE RV data, there is a signal at 6.9 days period but most likely related to the stellar rotation.
@@ 4.2 SPIRou radial velocities
We have achieved with SPIRou an average RV precision per measurement of 2.3 m s-1 and a rms of 9.1 m s-1 in the full LBL time-series. We display in the left panels of Fig. 3, the SPIRou LBL RV measurements as a function of time and their GLS periodogram. The periodogram displays a peak at days with a FAPbootstrap level of 2.2%. No peaks are observed at the rotational period of the star nor its harmonics. Our SPIRou LBL RV measurements are listed in Table LABEL:GL410_SPIROU_RVtable in the Appendix. A summary of the properties of the LBL RV time-series is given in Table 6.
Due to the reflex motion of the Earth in its orbit (i.e. the BERV), observations acquired at different moments of the year have different velocity shifts with respect to the center of telluric absorption lines. If a star has an intrinsic average systemic velocity with respect to the Sun, the total velocity of the star at the time of the observations is given by
In the near-IR the telluric absorption lines are abundant, thus the stellar photospheric absorption lines are affected by tellurics. The effect of telluric absorption is in first instance to deform the line profiles. After telluric correction, residuals from atmospheric lines, at a much lower level than original, may remain, and even for a perfect correction, the noise is higher in the affected velocity channels. The moment when the spectra is most affected by tellurics is when the atmospheric absorption is at the core of the stellar photospheric lines, that is when is between 10 and 10 km s-1, i.e. when the lines of chemical species common to the Earth’s atmosphere and the stellar photosphere (e.g. H2O, CO) are on the top of each other, in particular, in regions where the optical depth of the atmosphere is the greatest.
We have color-coded in Fig. 3 the RV measurements based . We display in blue the observations obtained when km s-1 and, in red, the observations taken when km s-1. The telluric correction of APERO is good, the LBL algorithm is efficient filtering outliers, and the zero-point correction in time and BERV space further corrects residuals of telluric lines. However, as can be seen in Fig. 3, at the m s-1 precision level, the systematic noise of tellurics is evident. Fig. 3, shows that a large fraction of the outliers in the RV time-series is in red-color. The effect of telluric lines becomes evident when the RV-time-series is plotted as a function of (see Fig. 3). The systematics due to the telluric absorption can be seen by eye, as an up and down correlated-noise pattern.
@@
| data-set | LBL | |||
| Time-series | ||||
@@
If the RV time-series is filtered of the data-points taken at km s-1 (i.e., the red-points in Fig. 3), the rms of the time-series (see Table 6), improves from 9.1 m s-1 for the full time-series to 6.5 m s-1 for the filtered time-series. If the periodogram is made only with observations taken at km s-1, we not only recover the peak at 6.02 days detected in the full time-series, but its significance increases from an FAPbootstrap=0.022 in the full time-series to FAPanalytical=1.4 in the filtered time-series. The case of Gl 410 dramatically shows the effect that telluric residuals can have adding RV noise thus blurring a planet detection.
@@
While filtering data-points taken at km s-1 is a robust and fast way of finding promising planet candidate signals, besides the obvious fact that it takes out a large number data points (39 measurements), it has the limitation that it can boost the power at certain frequencies because specific and periodic times are missing in the time sampling. The RV systematics and correlated noise generated by telluric residuals and detector based glitches and imperfections can be corrected. The LBL framework produces a 2D data set, a time series for each of the thousands LBL line measurements. In this 2D data-set, data-driven techniques such as principal component analysis (PCA) can be employed to further filter, mitigate and correct systematic noise in the RV measurements.
@@
In the context of the SPIRou collaboration, two PCA-based LBL RV correcting algorithms have been developed: Wapiti (Ould-Elhkim et al., 2023) and a weighted PCA method (wPCA, Artigau et al. in prep). We refer the reader to those papers for details on each methodology. The main difference between these two approaches is that Wapiti performs a PCA in the stellar velocity space, whereas in the second method the PCA is applied in the detector pixel space (inspired on Cretignier et al., 2023). We have re-analyzed the LBL RV data of Gl 410 with both of these methods. In the central and right panels of Fig. 3, we display the RV measurements and GLS periodograms obtained with wPCA and Wapiti. In Table 6, we present the summary of the properties of the RV time-series.
@@
Fig. 3 illustrates how wPCA and Wapiti correct for the systematic noise at km s-1 . The rms of the full time-series decreases from 9.1 m s-1 in the raw LBL to 6.8 m s-1 with wPCA and 6.3 m s-1 with Wapiti. The rms of the points at km s-1 (red dots in Fig. 3) decreases from 14.3 m s-1 in the raw LBL, to a 7.1 m s-1 with wPCA and 8.3 m s-1 with Wapiti. We also observe that the peak at 6.02 days becomes stronger in the full time-series and that it reaches a FAPbootstrap level of 3.5 Furthermore in the wPCA RV data, a new peak at 18.76 days with an FAPbootstrap level of 1.6% becomes visible. A peak at 18.73 days is also present in the Wapiti time-series, however with a negligible significance (FAPbootstrap=80%).
@@
In summary, the SPIRou RV data shows a clear periodic signal at a 6.02 days. The signal is not a harmonic or alias of the rotation period. There is evidence for a second tentative candidate signal at 18.7 days.
4.3 A test of robustness of the 6 day SPIRou signal.
@@
No periodicities are detected at 6.02 nor 18.7 days in the activity indicators time-series of SOPHIE nor SPIRou (details are provided in Sect. C in the Appendix). We performed an additional test of robustness to exclude that the signal at 6 days is due to activity. If a candidate signal is due to a planet (and not activity), it is expected that as more RV data points are accumulated the significance of the detection increases, and this should not depend on the way how the time-sampling of the measurements is performed.
@@
Using the km s-1 SPIRou wPCA time-series, we performed a test that consisted in checking the evolution of the FAP level of the 6 day peak in 100 different randomly sampled time-series. The construction of a RV time-series, starts by selecting randomly 10 measurements, calculating the GLS periodogram and measuring the FAP of the peak at 6 days. Then one randomly selected measurement is included in the time-series, the periodogram recalculated and the FAP of the peak at 6 days measured. The procedure is repeated until the complete number of visits is reached for a time-series. In this way, we constructed 100 independent evolutions of the FAP as a function of the number of measurements. This test implies to calculate around twelve thousand FAPs, to keep the computation time short, we calculated the FAP level using the analytical approximation implemented in pyAstronomy (Czesla et al., 2019). The analytical FAPs are a factor of a few lower than the FAPs calculated using a bootstrap method.
@@
The result of the test is displayed in Fig. 4. We can see that for all the time-series, the FAP level of the 6 day peak decreases as the number of measurements increases. In Fig. 4, we added the plot of the FAP level in the chronological time-series and the median FAP for each bin of number of measurements. The test indicates that the peak at =6 days in the periodogram is physical, and that it does not depend on the sampling of the time-series. This test gives confidence that the 6 day peak in the periodogram is indeed of Keplerian origin.
@@
5 Stellar activity filtering of the SOPHIE RVs and search for the Gl 410b signal.
@@
As opposed to a signal due to activity, a Keplerian reflex motion caused by the presence of a planet is achromatic. There is a hint of the 6 day period signal in the SOPHIE RV data of Gl 410. If the SOPHIE and SPIRou RVs are plotted phase-folded, both data-sets are consistent. However, the SOPHIE RVs have a much larger dispersion (14 m s-1) than the SPIRou RVs (7 m s-1). One interesting question is whether Gl 410b’s RV periodic signal could be retrieved in the SOPHIE data-set if the RV jitter induced by stellar activity is corrected.
@@
We tested two methods to search for the 6 day signal within the SOPHIE RV data. In the first method, the SOPHIE RVs time-series is analyzed with a model which is the sum of a quasi-periodic Gaussian Process (GP) and a Keplerian model. In the second method, we use the measurements of the longitudinal magnetic field () from SPIRou to train a GP that is used to correct the SOPHIE RVs for activity jitter, and then the planet signal is searched in the RV residuals.
@@
5.1 Method 1: SOPHIE GP modeling using the SOPHIE RV data alone.
@@
As the rotational period of Gl 410 is well constrained by spectropolarimetry, the SOPHIE RV data can be analyzed using a GP + Keplerian model. The goal here is to determine what model is more likely between a model without planet (i.e. only a GP) or a model with GP + planet. To perform the analysis we used RadVel (Fulton et al., 2018), and assumed a circular orbit. A Gaussian prior was used for the GP period with center at 13.9 days and =0.1 days corresponding to the stellar rotation period derived from the measurements of the longitudinal magnetic field (Donati et al., 2023). Uniform priors were set for: the planet period (0 to 20 days); the planet semi-amplitude (0 to 20 m s-1); the time of inferior conjunction (2459648.0 10 days); the GP amplitude (0 to 75 m s-1); the GP decay-time (0 to 2000 days); the GP smoothing-length (0 to 2 days); and the uncorrelated noise (0.0 to 10 m s-1).
@@
Table 4 provides a summary of the AIC tests comparing the models. The model with GP model alone (i.e. without planet) is statistically “ruled-out” by RadVel. In Table 13 in the Appendix, we present the posterior distributions obtained for the the model777 The stellar rotation period is well constrained from spectropolarimetry and provides a stringent prior on the cycle length of the GP which is recovered in the posterior distribution. We tested two models with a broader Gaussian distribution prior for the cycle length (13.93.5 and 13.97 days). All models converged to a GP cycle length of 13.840.03 days. . The Keplerian signal at 6.02 days is recovered. This test provides independent evidence for the existence of the 6 days planet inside the SOPHIE RV data. In Sect. 6.3, we will discuss the coherence of the Keplerian signals obtained with SPIRou and SOPHIE.
| Model | RMS | BIC | AICc | AICs | |
@@
5.2 Method 2: SOPHIE GP modeling using SPIRou’s as activity proxy.
@@
In this method, the search of the signal is done using a two step approach: in step one, a quasi-periodic Gaussian process (GP) model is fit to the SPIRou’s longitudinal magnetic field () time-series, and in step two, the posterior distribution of the hyper parameters (period, decay time and smoothing length) of the GP are used as prior for a quasi-periodic GP model of the SOPHIE RVs (other hyperparamers, ln , ln , priors are taken as uniform). The aim of this method is to filter the SOPHIE RVs for activity using as a proxy of the activity signal and search for the signal at 6 days in the residuals.
@@
We used the SPIRou measurements from Donati et al. (2023). In Sect. D in the Appendix, we describe the the quasi-periodic GP model used. In Fig. 22, also in the Appendix, we display the MCMC corner plots of the GP. We retrieve in our GP model of a stellar rotational period of days which is consistent with the previous determinations of the rotational period by Donati et al. (2023) and Fouqué et al. (2023). We modeled the SOPHIE RVs with the same quasi-periodic GP model and implementation used for . We focused on the SOPHIE RVs from 20212023 as this period has the quasi-simultaneous SPIRou measurements888Bellotti (2023) investigated in detail the SPIRou measurements of Gl 410 using GPs. One of their conclusions was that the decay time differed in the 2020/2021 and 2021/2022 data sets. Using ZDI, Bellotti (2023) found that the surface differential rotation was stronger in 2021/2022 inducing a faster surface variability of activity tracers. We use here a single GP for the complete time-series. Our GP, therefore, represent an ”average” that does not capture these fluctuating activity levels.. The prior and posterior distributions for the RV GP are given in Table 13 in the Appendix. In Fig. 23, also in the Appendix, we show the MCMC corner plots of the RV GP model. Fig. 5 summarizes the model results. The upper panel shows the GP model (in blue) and the SOPHIE RVs (black points). The middle panel displays the GP corrected SOPHIE RVs. The lower panel shows the GLS periodogram of the GP corrected SOPHIE RVs. After correcting for the activity jitter, the RV rms decreases from 12.7 m s-1 to 6.0 m s-1 and the GLS periodogram show a clear and strong peak at days with a FAPbootstrap of 2.4. This period is the same as that obtained in the SPIRou RV measurements.
@@
In Fig. 6, we display the phase-folded SPIRou RV time-series together with the SOPHIE GP corrected RVs, overlaying a Keplerian model fit with m s-1 and days. The data of both instruments are compatible. This provide us, one more time, further arguments to support the idea that the periodic signal at 6 days detected with SPIRou indeed has a Keplerian origin.
@@
We calculated a periodogram for the combined SPIRou (wPCA) and SOPHIE-GP corrected RV time-series. It is shown in Fig. 7. The periodogram of the combined SPIRou + SOPHIE-GP data-set (black lines) is very similar to that of the wPCA SPIRou data alone (green lines). However, the addition of the SOPHIE-GP RVs makes the FAP levels of the periodogram to decrease to lower powers (note the difference between the horizontal grey and green lines in the plot). The significance of the peak at days increases from a FAPanalytical=3.5 in the wPCA SPIRou data alone to a FAPanalytical=1.5 in the combined near-IR + optical data-set.

### v2

@@ 4 Radial velocity analysis
For the RV analysis, we used Generalized Lomb-Scargle (GLS) periodograms (Zechmeister & Kürster, 2009). To determine the false alarm probability levels (FAP), we used two methods. For peaks with a significance lower than 10-5 and the 10%, 1%, and 0.1% FAP levels in the periodograms, we used a bootstrap algorithm with 105 iterations. The method consists in generating a time series with dates equal to the real series of measurements, but assigning to each of them a RV measurement chosen randomly from the original data series. We labeled these FAP levels FAPbootstrap. For peaks with a significance higher than 10-5 (i.e., inferior to 1 / number of boostrap iterations), we used the GLS FAP analytical implementation of the package pyAstronomy (Czesla et al., 2019). We labeled these FAP levels FAPanalytical.
@@ 4.1 SOPHIE radial velocities
We achieved with SOPHIE an average RV precision per measurement of 2.4 m s-1 and an rms of 13.9 m s-1 in the full time series. Figure 2 displays the GLS periodograms obtained for the full data set, the 20102012, and 20212023 periods, respectively. Table 4 summarizes the properties of the time series and the strongest peak in the periodogram.
The full data set displays a peak at days with an FAPbootstrap of 4.0. This peak corresponds to the /2 harmonic of the stellar rotation period of days (Donati et al., 2023). Apart from the 1 day alias, there are no other peaks in the full data set periodogram. No signal is observed at the stellar rotation period. The 20102012 period shows no significant periodicities. A broad peak is observed close to days but its significance is low (FAP). In the 20212023 period, a peak at days appears at an FAPbootstrap level of 3.810-3. Except for the 1 day alias, no peaks are present in the periodogram with a significance below an FAPbootstrap=10%.
In summary, in the raw SOPHIE RV data, there is a signal at days, but it is most likely related to the stellar rotation.
| DATE | RJD | RV | RV | d2v | d3v | dLW | |||||||
@@ 4.2 SPIRou radial velocities
With SPIRou, we achieved an average RV precision per measurement of 2.3 m s-1 and an rms of 9.1 m s-1 in the full LBL time series. We display in the left panels of Fig. 3, the SPIRou LBL RV measurements as a function of time and their GLS periodogram. The periodogram displays a peak at days with a FAPbootstrap level of 2.2%. No peaks are observed at the rotational period of the star nor its harmonics. The SPIRou LBL RV measurements are listed in Table 9 (the full table is available at the CDS). A summary of the properties of the LBL RV time series is given in Table 12.
Due to the reflex motion of the Earth in its orbit (i.e., the BERV), observations acquired at different moments of the year have different velocity shifts with respect to the center of telluric absorption lines. If a star has an intrinsic average systemic velocity with respect to the Sun, the total velocity of the star at the time of the observations is given by
In the near-IR, the telluric absorption lines are abundant, thus the stellar photospheric absorption lines are affected by tellurics. The first effect of telluric absorption is the deformation of the photospheric line profiles. After telluric correction, residuals from atmospheric lines, at a much lower level than original, may remain, and even for a perfect correction, the noise is higher in the affected velocity channels. The moment the stellar spectra is most affected by tellurics is when the atmospheric absorption is at the core of the stellar photospheric lines, that is when is between 10 and 10 km s-1. In other words, this is when the absorption lines of chemical species common to the Earth’s atmosphere and the stellar photosphere (e.g., H2O, CO) are on the top of each other, particularly in regions where the optical depth of the atmosphere is the greatest.
We have color-coded in Fig. 3 the RV measurements based . We display the observations obtained when km s-1 in blue and the observations taken when km s-1 in red. The telluric correction of APERO is good, the LBL algorithm is efficient at filtering outliers, and the zero-point correction in time and BERV space further corrects residuals of telluric lines. However, as shown in Fig. 3, at the meter per second RV precision level, the systematic noise of tellurics is evident. Fig. 3 shows that a large fraction of the outliers in the RV time series are in red. The effect of telluric lines becomes evident when the RV time series is plotted as a function of (see Fig. 3). The systematics due to the telluric absorption can be seen, as an up and down correlated-noise pattern.
@@
| data set | LBL | |||
| Time series | ||||
@@
If the RV time series is filtered from the data-points taken at km s-1 (i.e., the red points in Fig. 3), the rms of the time series (see Table 12), improves: from 9.1 m s-1 for the full time series to 6.5 m s-1 for the filtered time series. If the periodogram is made only with observations taken at km s-1, we not only recover the peak at 6.02 days detected in the full time series, but its significance increases from an FAPbootstrap=0.022 in the full time series to FAPanalytical=1.4 in the filtered time series. The case of Gl 410 dramatically shows the effect that telluric residuals can have adding RV noise, ultimately blurring a planet detection.
@@
Filtering data-points taken at km s-1 is a robust and fast way of finding promising planet candidate signals. However, apart the obvious fact that this approach takes out a large of number data points (39 measurements), it also has the limitation that it can boost the power at certain frequencies because specific and periodic times are missing in the time sampling. The RV systematics and correlated noise generated both by telluric residuals and detector-based glitches and imperfections can be corrected. The LBL framework produces a 2D data set, a time series for each of the thousands of LBL line measurements. In this 2D data set, data-driven techniques such as principal component analysis (PCA) can be employed to further filter, mitigate, and correct systematic noise in the RV measurements.
@@
In the context of the SPIRou collaboration, two PCA-based LBL RV correcting algorithms have been developed: Wapiti (Ould-Elhkim et al., 2023) and a weighted PCA method (wPCA; Artigau et al. in prep). We refer to those papers for details on each methodology. The main difference between these two approaches is that Wapiti performs a PCA in the stellar velocity space, whereas in the second method, the PCA is applied in the detector pixel space (inspired on Cretignier et al., 2023). The Wapiti and wPCA RV measurements are given in Tables 9 and 11, respectively (both tables are available at the CDS). We re-analyzed the RV data of Gl 410 with both of these methods. In the central and right panels of Fig. 3, we display the RV measurements and GLS periodograms obtained with wPCA and Wapiti. In Table 12, we present the summary of the properties of the RV time series.
@@
Figure 3 illustrates how wPCA and Wapiti correct for the systematic noise at km s-1 . The rms of the full time series decreases from 9.1 m s-1 in the raw LBL to 6.8 m s-1 with wPCA and 6.3 m s-1 with Wapiti. The rms of the points at km s-1 (red dots in Fig. 3) decreases from 14.3 m s-1 in the raw LBL, to 7.1 m s-1 with wPCA and 8.3 m s-1 with Wapiti. We also observed that the peak at days becomes stronger in the full time series and that it reaches an FAPbootstrap level of 3.5. Furthermore in the wPCA RV data, a new peak at days with an FAPbootstrap level of 1.6% becomes visible. A peak at 18.73 days is also present in the Wapiti time series, however, it has a negligible significance (FAPbootstrap=80%).
@@
In summary, the SPIRou RV data shows a clear periodic signal at a days. The signal is not a harmonic or alias of the rotation period. There is evidence for a second, tentative, candidate signal at days.
4.3 A test of robustness of the 6-day SPIRou signal
@@
No periodicities were detected at 6.02 nor at 18.7 days in the activity indicators time series of SOPHIE and SPIRou (details are provided in Sect. B in the Appendix). We performed an additional test of robustness to exclude the possibility that the signal at 6 days is due to activity. If a candidate signal is due to a planet (and not activity), it is expected that as more RV data points are accumulated the significance of the detection increases, and this should not depend on the way the time-sampling of the measurements is performed.
@@
Using the km s-1 SPIRou wPCA time series, we performed a test that consists in checking the evolution of the FAP level of the days peak in 100 different randomly sampled time series. The construction of an RV time series starts by randomly selecting ten measurements, calculating the GLS periodogram, and measuring the FAP of the peak at days. One randomly selected measurement is then included in the time series, the periodogram is recalculated, and the FAP of the peak at days is measured. The procedure is repeated until the complete number of visits is reached for a time series. In this way, we constructed 100 independent evolutions of the FAP as a function of the number of measurements. This test implies the calculation of around twelve thousand FAPs. To keep the computation time short, we calculated the FAP level using the analytical approximation implemented in pyAstronomy (Czesla et al., 2019). The analytical FAPs are a factor of a few lower than the FAPs calculated using a bootstrap method.
@@
The result of the test is displayed in Fig. 4. One can see that for all the time series, the FAP level of the days peak decreases as the number of measurements increases. In Fig. 4, we include the plot of the FAP level in the chronological time series and the median FAP for each bin of number of measurements. The test indicates that the peak at days in the periodogram is physical and that it does not depend on the sampling of the time series. This test supports the argument that the days peak in the periodogram is indeed of Keplerian origin.
@@
5 Stellar activity filtering of the SOPHIE radial velocities and search for the Gl 410b signal
@@
As opposed to a signal due to activity, a Keplerian reflex motion caused by the presence of a planet is achromatic. There is a hint of the 6-day period signal in the SOPHIE RV data of Gl 410. If the SOPHIE and SPIRou RVs are plotted phase-folded, both data sets are consistent. However, the SOPHIE RVs have a much larger dispersion (14 m s-1) than the SPIRou RVs (7 m s-1). One interesting question is whether Gl 410b’s RV periodic signal could be retrieved in the SOPHIE data set if the RV jitter induced by stellar activity is corrected.
@@
We tested two methods to search for the 6-day signal within the SOPHIE RV data. In the first method, the SOPHIE RVs time series is analyzed with a model that is the sum of a quasi-periodic GP and a Keplerian model. In the second method, we used the measurements of the longitudinal magnetic field () from SPIRou to inform a GP that is used to correct the SOPHIE RVs for activity jitter, and then we searched for the planet signal in the RV residuals.
@@
5.1 Method 1: SOPHIE Gaussian process modeling using the SOPHIE radial velocity data alone
@@
As the rotational period of Gl 410 is well constrained by spectropolarimetry, the SOPHIE RV data can be analyzed using a GP plus Keplerian model. The goal here is to determine what model is more probable between a model without a planet (i.e., only a GP) and a model with a GP plus a planet. To perform the analysis, we used RadVel (Fulton et al., 2018) and assumed a circular orbit. A Gaussian prior was used for the GP period, with center at 13.9 days and =0.1 days, which corresponds to the stellar rotation period derived from the measurements of the longitudinal magnetic field (Donati et al., 2023). Uniform priors were set for the planet period (0 to 20 days), the planet semi-amplitude (0 to 20 m s-1), the time of inferior conjunction (2459648.0 10 days), the GP amplitude (0 to 75 m s-1), the GP decay time (0 to 2000 days), the GP smoothing length (0 to 2 days), and the uncorrelated noise (0.0 to 10 m s-1).
@@
Table 7 provides a summary of the small sample Akaike information criterion (AICc) tests comparing the models. The model with a GP model alone (i.e., without planet) was statistically “ruled out” by RadVel. In Table 13 in the Appendix, we present the posterior distributions obtained for the model131313The stellar rotation period is well constrained from spectropolarimetry and provides a stringent prior on the cycle length of the GP, which is recovered, as expected, in the posterior distribution. We tested two models with a broader Gaussian distribution prior for the cycle length (13.93.5 and 13.97 days). All models converged to a GP cycle length of 13.840.03 days. . The Keplerian signal at days is retrieved in the model. This test provides independent evidence for the existence of the 6-day planet inside the SOPHIE RV data. In Sect. 6.3, we discuss the coherence of the Keplerian signals obtained with SPIRou and SOPHIE.
| Model | RMS | BIC | AICc | AICc | |
@@
5.2 Method 2: SOPHIE Gaussian process modeling using SPIRou’s as an activity proxy
@@
In this second method, the search of the signal is done using a two step approach. In step one, a quasi-periodic GP model is fit to the SPIRou’s longitudinal magnetic field () time series, and in step two, the posterior distribution of the hyper parameters (period, decay time, and smoothing length) of the GP are used as prior for a quasi-periodic GP model of the SOPHIE RVs (other hyperparamers, such as ln , and ln , priors are taken as uniform). The aim of this method is to filter the SOPHIE RVs for activity using as a proxy of the activity signal and search for the signal at 6 days in the residuals.
@@
We used the SPIRou measurements from Donati et al. (2023). In Sect. C in the Appendix, we describe the quasi-periodic GP model we used. In Fig. 22, also in the Appendix, we display the MCMC corner plots of the GP. We retrieved in our GP model of a stellar rotational period of days, which is consistent with the previous determinations of the rotational period by Donati et al. (2023) and Fouqué et al. (2023). We modeled the SOPHIE RVs with the same quasi-periodic GP model and implementation used for . We focused on the SOPHIE RVs from 2021 to 2023, as this period has the quasi-simultaneous SPIRou measurements141414Bellotti (2023) investigated in detail the SPIRou measurements of Gl 410 using GPs. One of their conclusions was that the decay time differed in the 20202021 and 20212022 data sets. Using ZDI, Bellotti (2023) found that the surface differential rotation was stronger in 2021 to 2022 inducing a faster surface variability of activity tracers. Here, we use a single GP for the complete time series. Therefore, our GP represents an “average” that does not capture these fluctuating activity levels.. The prior and posterior distributions for the RV GP are given in Table 13 in the Appendix. In Figure 23, also in the Appendix, we show the MCMC corner plots of the RV GP model. Figure 5 summarizes the model results. The upper panel shows the GP model (in blue) and the SOPHIE RVs (black points). The middle panel displays the GP-corrected SOPHIE RVs. The lower panel shows the GLS periodogram of the GP-corrected SOPHIE RVs. After correcting for the activity jitter, the RV rms decreases from 12.7 m s-1 to 6.0 m s-1, and the GLS periodogram shows a clear and strong peak at days with a FAPbootstrap of 2.4. This period is the same as that obtained in the SPIRou RV measurements.
@@
In Figure 6, we display the phase-folded SPIRou RV time series together with the SOPHIE GP-corrected RVs, while overlaying a Keplerian model fit with m s-1 and days. The data of both instruments are compatible. This, provided us, one more time, with further arguments to support the idea that the periodic signal at 6 days detected with SPIRou indeed has a Keplerian origin.
@@
We calculated a periodogram for the combined SPIRou (wPCA) and SOPHIE minus GP RV time series. It is shown in Fig. 7. The periodogram of the combined SPIRou and SOPHIE minus GP data set (black lines) is very similar to that of the wPCA SPIRou data alone (green lines). However, the addition of the SOPHIE minus GP RVs makes the FAP levels of the periodogram decrease to lower powers (we emphasize the difference between the horizontal gray and green lines in the plot). The significance of the peak at days increases from an FAPanalytical=3.5 in the wPCA SPIRou data alone to an FAPanalytical=1.5 in the combined near-IR plus optical data set.